Harry and Larry are a married couple in California. They itemize deductions. Before making a charitable contribution in 2018 their taxable income is $182,000. Their average tax rate is 20% and their effective tax rate is 18%.

If they decide to make a deductible charitable contribution of $10,000 their federal income tax for 2018 will decrease by $2,800.

True or False
Answer #1

False

Taxable income before any charitable deduction is $182000

So tax on 182000= 182000*18%=$32760

If they decide to make deduction of charitable contribution

There income shall be 182000-10000=$172000

And tax shall be 172000*18%=$30960

Tax will decrease by 32760-30960=$1800